Suppose 100 g of PbS is reacted with 4g of C and 20g O2.

Find the limiting reactant for this reaction.   

1
Expert's answer
2021-11-11T10:29:03-0500

2PbS(s) + 2C(s) + 3O2(g) → 2Pb(s) +2CO(g) + 2SO2(g)

n(PbS) = 100 g / (2 *239.3 g/mol) = 0.21 mol

n(C) = 4 g / (2*12 g/mol) = 0.17 mol

n(O2) = 20 g / (3*32 g/mol) = 0.21 mol

C is limiting reactant